Mathematics 133 – Intensive Calculus for Science 1 “Make-up Quiz 0” – Algebra Review – September 6, 2005

Your Name:

Do all work on this sheet (front and back). You have 20 minutes. No calculators on this quiz.

A) Write as a single quotient and simplify: \frac{1}{2+5x^{2}} + \frac{1+x^{2}}{1−x^{2}}

B) Solve for u: 4(2 − 3u) + 1 = 7 − 2u.

C) Solve for x: 3x^{2} + 3 = 10x (find all solutions)

D) Solve for t: t^{2/3} = 8t E) Simplify: (x^{4/3})^{5} · x^{−2/3} F) Expand and simplify: (a − b + 1)(a + b + 1) G) Find the coordinates of the center and the radius of the circle with equation x^{2} +

16x + y^{2} − 2y = 0.